There are pigs and chicken in the farm. The total no. of heads and legs of both animals is 100. How many pigs and chicken are there in the farm ?

Answer by ankor@dixie-net.com(22740)

There are pigs and chicken in the farm. The total no. of heads and legs of both animals is 100. How many pigs and chicken are there in the farm?
:
Let p = no. of pigs
Let c = no. of chicks
:
Legs expression:
4p + 2c
;
Head expression:
p + c
:
no. of legs = 100 - no. of heads
:
(4p + 2c) = 100 - (p+c)
4p + 2c = 100 - p - c
4p + p + 2c + c = 100
5p + 3c = 100
5p = 100 - 3c
p = 100/5 - (3/5)c
p = 20 - .6c
;
p has to be an integer, what integer value for c will provide this?
:
Turns out c = multiples of 5 from 5 thru 30 will satisfy
c p
5 17
10 14
15 11
20 8
25 5
30 2
:
Examples
5 chicks have 10 legs + 5 heads total = 15
17 pigs have 68 legs + 17 heads total = 85
and
30 chicks have 60 legs + 30 heads total = 90
2 pigs have 8 legs + 2 heads total = 10